OpenStudy (anonymous):

Please help!!!!!!! Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Which one represent sale price in here?

OpenStudy (anonymous):

@lalaly

OpenStudy (calculusxy):

P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (anonymous):

is it 40%?

OpenStudy (calculusxy):

Since it is a profit of 40% you just add a whole to it making it 1.4%

OpenStudy (anonymous):

I don't want it to solve, just want to know which one is sale price

OpenStudy (anonymous):

is it 40%?

OpenStudy (calculusxy):

Yes

OpenStudy (anonymous):

thank you so much for help!!!

OpenStudy (calculusxy):

Welcome :)

OpenStudy (anonymous):

@tkhunny @hartnn I wanted to know, 40% is the one that represented sale price here. Is it true?

OpenStudy (anonymous):

@ybarrap

OpenStudy (tkhunny):

"They set the sale price of each flower " This is bad. It should have said "sell price", not "sale price". A "sale" is normally a discount, and that's just not what they are talking about.

OpenStudy (anonymous):

So it is 40% ?

OpenStudy (anonymous):

I know they are not talking about that. I just wanted to know it

OpenStudy (tkhunny):

The SELL price is $0.50 + 40%($0.50) = $0.50 + $0.20 = $0.70 The Profit on each flower is that markup. 40%($0.50) = $0.20.

OpenStudy (anonymous):

so it is not called sale price?

OpenStudy (anonymous):

40% is a mark up

OpenStudy (anonymous):

@Compassionate

OpenStudy (tkhunny):

Absolutely not talking about a "sale" price, here. It might mean the price AT THE SALE, I suppose, but that is confusing. You can have a "sale" and "sell" things, but when talking about price, the "sale" price is a discount off the regular retail.

OpenStudy (tkhunny):

You must be VERY careful when talking about "markup" There are two ways to calculate it. In this case, it is a percentage (40%) of the Wholesale price. Sometimes, and you must read carefully to know, it is a percentage of the Retail Price. That is a different animal than what we have, here.

OpenStudy (anonymous):

So what exactly represent 40 percent here?

OpenStudy (tkhunny):

40% of the Wholesale price is the Markup in this arrangement.

OpenStudy (anonymous):

ok thank a lot for help!!!!! And I am sorry for trouble. :(

OpenStudy (tkhunny):

What trouble? It's why we're here. :-)